A chemist has 10% and 16% acid solutions. How many liters of each solution are needed to produce 30 liters of a 14% acid solution?
Answer by CharlesG2(834)

A chemist has 10% and 16% acid solutions. How many liters of each solution are needed to produce 30 liters of a 14% acid solution?
% in A * amount A + % in B * amount B = % in C * amount C
10% * x + 16% * (30 - x) = 14% * 30
0.10x + 0.16(30 - x) = 0.14 * 30
0.10x + 4.8 - 0.16x = 4.2
-0.06x = -0.6 (rearranged)
0.06x = 0.6 (divided both sides by -1)
x = 10
10 liters of 10% and 20 liters of 16%